OpenStudy (anonymous):

Sue drove her car at 25 mph for the first part of the trip and then changed her speed to 75 mph for the second part. If the entire trip is 725 miles and took 13 hours, how many hours are traveled at each speed?

OpenStudy (anonymous):

first what is the formula in finding speed of the car..?

OpenStudy (anonymous):

idk let me see

OpenStudy (anonymous):

x+13=725?

OpenStudy (anonymous):

time = distance / speed so, 13 = x/25 + (725-x)/75 where x is distance of the first part of the trip

OpenStudy (anonymous):

125?

OpenStudy (anonymous):

x = 125. use the formula above to find the time for each part of the trip

OpenStudy (anonymous):

13=5+600/75?

OpenStudy (anonymous):

it's bout 10 hours at 25 and 3 at 75 because it's 3x faster?

OpenStudy (anonymous):

is this right?